OpenStudy (anonymous):

Can someone please please help me with some more surface area of pyramids and cones???

OpenStudy (anonymous):

maybe i might be able to help little bit maybe

OpenStudy (anonymous):

take a screen shot of problem and upload it

OpenStudy (anonymous):

for this one i need to find the slant hight, lateral area, and the surface area

OpenStudy (anonymous):

for this one i need to find the slant height

OpenStudy (anonymous):

okay hold on sec

OpenStudy (anonymous):

okay on the slant height you use this formula =Lateral Area of a regular pyramid: ½ perimeter × slant height = 96 on this one

OpenStudy (anonymous):

because see tha base of it is got dotted line which means you take half of 16 = 8 and * that by 12 which gives us 96

OpenStudy (anonymous):

on frist one to get surface area use this formula Surface Area = Lateral Area + n × Bases.

OpenStudy (anonymous):

i should of put this at the top... they are multiple choice.. for the cone the choices are a) 20 b) 10.6 C08.9 d) 14.4 e)7.1

OpenStudy (anonymous):

well hummmmmmmm

OpenStudy (anonymous):

yeah sorry im trying to do like 5 things at once

OpenStudy (anonymous):

all good i am trying to do my math work and help others

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

(25/4+144/4)^(1/2) = (169/4)^(1/2) = 13/2 (169/4+25/4)^(1/2) = 6.96 for side of pyramid

OpenStudy (anonymous):

6.96 for side of pyramid so would this be the slant height?

OpenStudy (anonymous):

?, the side going from one of the bottom corners to the top.

OpenStudy (anonymous):

It's the longest side on the pyramid

OpenStudy (anonymous):

for the pyramid: the slant height choices are: a)4.5cm b)5.5cm c)6.5cm d)6cm e)7.8 for the lateral area choices are: a)57.5cm^2 b)65cm^2 c)32.5cm^2 d)78cm^2 e)90cm^2 for the surface area: a)57.5cm^2 b)65cm^2 c)32.5cm^2 d)78cm^2 e)90cm^2

OpenStudy (anonymous):

Ok, slant height is the height from the midpoint on the base to the top. So 6.5

OpenStudy (anonymous):

for SA, .5bh*4 + b^2 b=5, h=6.5 (5)(6.5)(2)+5(5) 90cm squared

OpenStudy (anonymous):

I'm guessing lateral area means without the base, so it's probably 65

OpenStudy (anonymous):

All you need to know to do these problems to use the Pythagorean theorem on Right triangles to find the missing side, and area of triangles is simply .5base times height.

OpenStudy (anonymous):